# TAR3D: Creating High-Quality 3D Assets via Next-Part Prediction

### v1

@@ Abstract
We present TAR3D, a novel framework that consists of a 3D-aware Vector Quantized-Variational AutoEncoder (VQ-VAE) and a Generative Pre-trained Transformer (GPT) to generate high-quality 3D assets. The core insight of this work is to migrate the multimodal unification and promising learning capabilities of the next-token prediction paradigm to conditional 3D object generation. To achieve this, the 3D VQ-VAE first encodes a wide range of 3D shapes into a compact triplane latent space and utilizes a set of discrete representations from a trainable codebook to reconstruct fine-grained geometries under the supervision of query point occupancy. Then, the 3D GPT, equipped with a custom triplane position embedding called TriPE, predicts the codebook index sequence with prefilling prompt tokens in an autoregressive manner so that the composition of 3D geometries can be modeled part by part. Extensive experiments on ShapeNet and Objaverse demonstrate that TAR3D can achieve superior generation quality over existing methods in text-to-3D and image-to-3D tasks.
Figure 1: Examples of high-quality 3D asset gallery generated by our TAR3D model.
@@ 1 Introduction
Conditional 3D object generation aims to generate high-quality 3D assets that semantically conform to the given prompts, e.g., 2D images, and text. It has witnessed significant progress with substantial advances in diffusion-based methods, including Score Distillation Sampling (SDS) optimization [46, 7, 71, 39], multiview synthesis [37, 52, 68, 26], and 3D-aware diffusion generation [75, 67, 30]. However, establishing a unified model across various modalities poses huge challenges due to the different paradigms of diffusion models and the models for other modalities.
On another line, multimodal large language models based on the next-token prediction paradigm, such as GPT-4 [1], Qwen [3], and InternVLs [12, 42], have revolutionized the field of language generation with groundbreaking reasoning abilities and incredible scalability. A series of representative works, e.g., LlamaGen [55], and Lumina-mGPT [34], also have been exploring the envelope of autoregressive models in image generation. These achievements demonstrate that the next-token-prediction scheme has the potential to become a promising path toward artificial general intelligence. More recently, pioneering works like MeshGPT [53] and MeshXL [8] attempt to introduce such an autoregressive manner into 3D mesh generation. However, they usually quantize the vertexes of mesh faces into discrete embeddings, resulting in a sequence whose length is 9 times the number of faces. When handling high-quality 3D assets, the excessively long sequence will substantially burden the modeling process and GPU memory.
@@
In contrast to the above methods, we propose a triplane-driven novel TAR3D framework, consisting of a 3D VQ-VAE and a 3D GPT. Particularly, the 3D VQ-VAE encodes the 3D shapes into compact triplane latent features with fixed height and width, from which a trainable codebook of context-rich 3D geometry parts is employed to acquire a set of discrete embeddings. Such a manner allows the 3D meshes to be represented with a fixed-length sequence, regardless of the number of faces, reducing the reliance on huge GPU resources. After that, these quantized latent representations are decoded to a neural occupancy field for 3D reconstruction. We empirically observe that the reconstructed 3D meshes of such naive VQ-VAE are prone to losing fine-grained geometry details. To address this limitation, we propose to enhance the feature map of each plane and promote information exchange among different planes by incorporating feature deformation and attention mechanism designs during the decoding process. The 3D GPT is driven by prefilling prompt embeddings to model the codebook index sequence corresponding to quantized triplane features, achieving conditional 3D object generation in an autoregressive manner. To preserve as much spatial information as possible, we custom-craft a 3D positional encoding strategy termed TriPE for this sequence, in which the 2D information of each plane and the 1D ones between the same position of the three planes are organically fused.
@@
To validate the effectiveness of our TAR3D framework, we conduct extensive experiments on a wide range of 3D objects of two popular benchmark datasets, i.e., ShapeNet [6] and Objevaerse [15]. Based on the autoregressive manner and our well-designed strategies, the TAR3D can create high-quality 3D assets, as shown in Fig. 1. The quantitative and quantitative results also demonstrate that our TAR3D can significantly outperform recent cutting-edge conditional 3D generation methods.
@@
  • We present TAR3D, a novel autoregressive pipeline composed of 3D VAE and GPT for conditional 3D object generation. To the best of our knowledge, it is the first attempt to quantize the 3D objects with triplane representations and generate high-quality assets part by part.
@@ 2 Related Work
In recent years, research on 3D generation has seen rapid development. Early approaches mainly focus on the generation of different forms of 3D models, e.g., point clouds [66], meshes [9], and volumes [5] in a text/image-conditioned or unconditional manner. Limited by the categories and quantity of 3D objects used for training, the generalization ability of these methods is usually poor. With the remarkable progress achieved by diffusion models in 2D image generation, a large number of researchers have explored migrating the pre-trained 2D priors to 3D generation. Pioneer works like DreamFusion [46], SJC [61], and Fantasia3D [7] feed the rendered views to a pre-trained 2D diffusion model and perform per-shape optimization for knowledge distillation. Despite ushering in a new era, these methods suffer from a series of serious issues, e.g., time-consuming and multi-face. Another line of methods like Zero123 [37], Zero123++ [52], and One2345 [35] use the rendered views of 3D objects to finetune the pre-trained diffusion models for new-view or multi-view generation. To improve the consistency between the generated multiple views, Consistent123 [33] and Cascade-Zero123 [10] introduce extra priors, e.g., boundary, and redundant views. SyncDreamer [38] proposes to correlate the corresponding features across different views by building a 3D-aware attention mechanism. These methods can be followed by a sparse-view reconstruction model, e.g., NeRFs [44, 62, 45], LRMs [24, 68, 57], to generate 3D objects. However, such indirect generation fashion may lead to detail loss or reconstruction failures due to their heavy reliance on the fidelity of multi-view images.
@@ 2.2 VAE & VQ-VAE
VAE, short for Variational Autoencoder [29], consists of an encoder for compression and a decoder for reconstruction. It is usually utilized to map high-dimensional input information to continuous probabilistic latent representations and has a far-reaching impact in the field of generative modeling. Thanks to this work, plenty of diffusion models [23, 50, 51] can be trained on limited computational resources while retaining their quality and flexibility. Recent 3D generation approaches like Clay [75], Direct3D [67], and LN3Diff [30] also explore constructing 3D-aware diffusion models based on this technology. Vector Quantised-Variational AutoEncoder (VQ-VAE) [60] is a variant of VAE. It introduces the codebook mechanism to quantize the continuous latent representations into discrete components, achieving promising performance on many generative tasks, e.g., text-to-image generation [49], music generation [18, 17], and speech gesture generation [2]. VQ-GAN [19] proposes to promote VQ-VAE by incorporating adversarial loss during the training process.
In this paper, we build a 3D-aware VQ-VAE to quantize the triplane features of 3D shapes and acquire discrete geometric parts for autoregressive 3D generation.
@@ 2.3 GPT
Generative Pre-trained Transformer (GPT) [47, 43, 77, 78] originate in the field of natural language processing (NLP), with the core idea being to leverage large-scale unsupervised pretraining to learn the underlying distribution of language from massive text data. Based on the decoder-only transformer architecture, it autoregressively generates text sequences according to the next-token-prediction paradigm. This series of works [1, 80, 3, 12, 42] with groundbreaking reasoning ability and incredible scalability continue to emerge, revolutionizing language generation.
Inspired by these achievements, many researchers have attempted to transfer this scheme to image generation. For example, Parti [70] proposes a pathways autoregressive text-to-image model, which regards image generation as sequence-to-sequence modeling for high-fidelity images. VAR [58] redefines the autoregressive learning on images as coarse-to-fine “next-scale prediction”. LlamaGen [55] verifies that vanilla autoregressive models, e.g., Llama [59], can achieve state-of-the-art image generation performance without inductive biases on visual signals if scaling properly. Emu3 [64] achieves excellent performance in both generation and perception tasks by tokenizing images, text, and videos into a discrete space. More recently, several works like MeshGPT [53], MeshAnything [11], and MeshXL [8] also explore to generate the faces of 3D meshes autoregressively with GPT.
@@ 3 Methodology
3D shape encoder is designed to acquire compact and robust latent representations of 3D objects for fine-grained geometry information. Given a 3D object, we first uniformly sample high-resolution point clouds from its surface. To enhance the expressive power of the point clouds, the corresponding normals are also included in the point cloud representation that is denoted as $\mathrm{P}\in\mathbb{R}^{B\times N_{p}\times(3+3)}$, where $B$ is the batch size, and $N_{p}$ is the number of points. Then, we apply the Fourier positional  [56] encoding to the point cloud representation to capture high-frequency details. Inspired by previous works for point cloud understanding [27, 73], we employ a transformer-based architecture consisting of a cross-attention layer and $N_{1}$ self-attention layers to extract the latent features of the 3D point cloud. More precisely, the point cloud information is injected into a series of learnable query tokens that is denoted as $\mathrm{e}\in\mathbb{R}^{B\times(3\times h\times w)\times d_{e}}$, where $h$ and $w$ are the height and width of the triplane feature maps respectively, and $d$ represents the channel number of these learnable tokens, via the cross-attention layers. After that, the representational ability of these tokens is enhanced by the following self-attention layers, resulting in triplane latent representations, i.e., $\hat{\mathrm{z}}\in\mathbb{R}^{B\times(3\times h\times w)\times d_{z}}$.
@@ 4.1 Experiment Setup
Datasets. To examine the effectiveness of our TAR3D, we conduct experiments on two widely used open-source 3D datasets, i.e., ShapeNet [6] and Objaverse [15]. To be specific, the ShapeNet dataset provides 52,472 manufactured meshes covering 55 categories. We follow the splits from 3DILG [72], where 48,597 samples are used for training, 1,283 for validation, and 2,592 for testing. Inspired by previous works on data filtering [31, 75], we score the rendered normal maps of 800,000 meshes in the Objaverse dataset and obtain about 210,000 geometry objects. Moreover, 1,000 samples are randomly selected for performance evaluation and the remaining ones are employed for model training. For each 3D asset in these two datasets, we adopt the rendered images and textual descriptions from ULIP [69] to build the prompt system. In particular, we uniformly select 4 images from all the rendered images, whose top 1 captions are utilized as the text prompts.
Metrics. We evaluate the performance of the methods from two aspects: 2D visual quality and 3D geometric quality. For the 2D visual evaluation, we compare the novel views rendered from the synthesized 3D mesh with the ground truth views based on a set of common metrics, including Peak Signal-to-Noise Ratio (PSNR), Perceptual Loss (LPIPS) [76], Structural Similarity (SSIM) [65], and CLIP [48] score. For the 3D geometric evaluation, we compare the point clusters that are randomly sampled from the generated meshes and the ground-truth meshes. Following the protocol in previous works [68, 36], we employ the chamfer distance value and the F-Score with a threshold of 0.2 as the primary evaluation metrics.
Figure 4: Visual comparisons of the 3D meshes generated by our TAR3D and recent state-of-the-art approaches on image-to-3D object generation. Given the same input images, our TAR3D can produce 3D assets with superior geometric details against other baselines.
@@
In our 3D GPT, we adopt the pre-trained DINO [74](ViT-B16) and FLAN-T5 XL [14] to encode the conditional images and text prompts respectively. As far as the decoder-only transformer, we follow the GPT-L setting of LLamaGen [55], which consists of 24 transformer layers with a head number of 16 and a dimension of 1024. We employ the AdamW optimizer with a learning rate of $41e\mathrm{-}$ 4 and a total batch size of 80 to train our 3D GPT for 100K steps. In addition, the classifier-free guidance (CFG) [22] scale of 5 is also introduced in the autoregressive inference to improve geometry quality and image/text-3D alignment.
@@ 4.2 Qualitative Evaluation
We conduct visualization experiments on a wide range of methods for image-to-3D and text-to-3D tasks to highlight the effectiveness of the proposed TAR3D model. In particular, we select these methods included in the comparison based on three principles: a) recently published, b) representative, and c) with official open-source code.
Image-to-3D. We first provide visualization comparisons between our TAR3D with recent state-of-the-art models for image-to-3D generation, including two multiview-based methods like SyncDreamer [38], and InstantMesh [68], and two 3D diffusion-based methods like Shap-E [28] and Michelangelo [79]. As shown in the airplane and hammer samples of Fig. 4, the multi-view methods may produce 3D objects with discontinuous geometry parts or even incorrect objects, which are attributed to inconsistencies in their views used for reconstruction. Meanwhile, the motorcycle sample shows that diffusion-based methods are prone to generating noisy 3D objects, possibly due to anomalies in the conditional denoising process. In contrast to these approaches, our TAR3D, equipped with the powerful learning capabilities of GPT, can synthesize high-quality 3D objects with superior geometric details.
Figure 5: Given a wide range of input images at specific views, our TAR3D model is capable of inferring the full picture of the primary objects and producing high-quality 3D meshes that adhere to them.
@@
Shap-E [28] SyncDreamer [38] Michelangelo [79] InstantMesh [68] TAR3D (Ours)
PSNR $\uparrow$ 10.991 11.269 11.928 11.560 13.626
SSIM $\uparrow$ 0.702 0.706 0.734 0.721 0.763
Clip-Score $\uparrow$ 0.834 0.837 0.864 0.847 0.868
LPIPS $\downarrow$ 0.325 0.320 0.278 0.303 0.216
Chamfer Distance $\downarrow$ 0.156 0.158 0.117 0.137 0.066
F-Score $\uparrow$ 0.163 0.178 0.226 0.179 0.303

### v2

@@ Abstract
We present TAR3D, a novel framework that consists of a 3D-aware Vector Quantized-Variational AutoEncoder (VQ-VAE) and a Generative Pre-trained Transformer (GPT) to generate high-quality 3D assets. The core insight of this work is to migrate the multimodal unification and promising learning capabilities of the next-token prediction paradigm to conditional 3D object generation. To achieve this, the 3D VQ-VAE first encodes a wide range of 3D shapes into a compact triplane latent space and utilizes a set of discrete representations from a trainable codebook to reconstruct fine-grained geometries under the supervision of query point occupancy. Then, the 3D GPT, equipped with a custom triplane position embedding called TriPE, predicts the codebook index sequence with prefilling prompt tokens in an autoregressive manner so that the composition of 3D geometries can be modeled part by part. Extensive experiments on several public 3D datasets demonstrate that TAR3D can achieve superior generation quality over existing methods in text-to-3D and image-to-3D tasks.
Figure 1: 3D asset gallery generated by the proposed TAR3D model. We employ SyncMVD [40] as the texture synthesizer.
@@ 1 Introduction
Conditional 3D object generation aims to generate high-quality 3D assets that semantically conform to the given prompts, e.g., 2D images, and text. It has witnessed significant progress with substantial advances in diffusion-based methods, including Score Distillation Sampling (SDS) optimization [48, 7], multiview synthesis [38, 54, 70], and 3D-aware diffusion generation [76, 68, 32]. However, establishing a unified model across various modalities poses huge challenges due to the different paradigms between diffusion models and large language models (LLMs) [1, 3, 12, 43].
More recently, pioneering works like MeshGPT [55] and MeshXL [8] attempt to introduce the autoregressive manner of LLM into 3D mesh generation. However, they struggle in a sequence length of 9 times the face number, which is attributed to their quantization of the vertices of mesh faces. Such an excessively long sequence length limits their applicability in industrial 3D assets with hundreds of thousands of faces. This motivates us to investigate whether it is possible to encode a mesh with an arbitrary number of faces into a fixed-length sequence to alleviate the computational burden of autoregressive modeling.
To address this question, the key is to efficiently encode the 3D representations such that the sequence length is not proportional to the number of faces. Triplane representation provides such a good property. Different from other 3D representations like unordered point clouds or verbose voxels, this representation is able to balance storage efficiency with strong expressiveness by compressing 3D information into three 2D feature maps with a fixed size. Due to the high-level 2D compositions, it fits discretization better and can also take advantage of the strategies from image quantization methods, e.g., LLamagen [57]. Based on triplane representation, we propose TAR3D, a novel 3D autoregressive framework consisting of a 3D VQ-VAE and a 3D GPT.
@@
The 3D VQ-VAE encodes the 3D shapes into compact triplane features, from which a trainable codebook of context-rich 3D geometry parts is employed to acquire a set of discrete embeddings. In this way, the 3D meshes can be represented as a feature sequence with triplane-size length, regardless of the number of faces, reducing the reliance on huge GPU resources. The quantized latent representations are then decoded to a neural occupancy field for 3D reconstruction. To achieve information exchange among different planes, we propose incorporating feature deformation and attention mechanism designs during the decoding process for fine-grained geometry details. The 3D GPT is driven by prefilling prompt embeddings to model the codebook index sequence corresponding to quantized triplane features, enabling conditional 3D object generation in an autoregressive manner. To preserve more spatial information, during generation, we also custom-craft a 3D positional encoding strategy dubbed TriPE, in which the 2D positional information of each plane and the 1D ones between the same position of the three planes are organically fused.
@@
To validate the effectiveness of our TAR3D, we conduct extensive experiments on a wide range of 3D objects of two popular benchmark datasets, i.e., ShapeNet [6] and Objaverse [16], and an out-of-domain dataset, Google Scanned Objects [20]. Based on the autoregressive manner and our well-designed strategies,TAR3D can create high-quality 3D assets, as shown in Fig. 1. The quantitative and qualitative results also demonstrate that our TAR3D can significantly outperform recent cutting-edge 3D generation methods.
@@
  • We present TAR3D, a novel autoregressive pipeline composed of a 3D VAE and a 3D GPT for conditional 3D object generation. To our knowledge, this is the first attempt to quantize the 3D objects with triplane representations and generate high-quality assets part by part.
@@ 2 Related Work
Early 3D generation approaches mainly focus on the generation of different forms of 3D models, e.g., point clouds [67], meshes [9], and volumes [5] in a text/image-conditioned or unconditional manner. Limited by the categories and quantity of 3D objects used for training, these methods often do not generalize well. With the remarkable progress achieved by diffusion models in 2D image generation, a large number of researchers have explored migrating the pre-trained 2D priors to 3D generation. Pioneer works like DreamFusion [48], SJC [62], and Fantasia3D [7] feed the rendered views to a pre-trained 2D diffusion model and perform per-shape optimization for knowledge distillation. Despite ushering in a new era, these methods suffer from a series of serious issues, e.g., time-consuming and multi-face. Another line of methods like Zero123 [38], Zero123++ [54], and One2345 [36] use the rendered views of 3D objects to finetune the pre-trained diffusion models for new-view or multi-view generation. To improve the consistency between the generated multiple views, Consistent123 [35] and Cascade-Zero123 [10] introduce extra priors, e.g., boundary, and redundant views. SyncDreamer [39] proposes to correlate the corresponding features across different views by building a 3D-aware attention mechanism. These methods can be followed by a sparse-view reconstruction model, e.g., NeRFs [45, 63, 47], LRMs [27, 24, 70] and LGM [59], to generate 3D objects. However, the indirect generation fashion of these methods may lead to detail loss or reconstruction failures due to their heavy reliance on the fidelity of multi-view images.
@@ 2.2 VAE & VQ-VAE
Variational Autoencoder (VAE) [31] is usually utilized to map high-dimensional input information to continuous probabilistic latent representations and has a far-reaching impact in the field of generative modeling. Thanks to this work, recent diffusion models [26, 52, 53] can be trained on limited computational resources while retaining their quality and flexibility. Recent 3D generation approaches like Clay [76], Direct3D [68], and LN3Diff [32] also explore to construct 3D-aware diffusion models based on this technology. Vector Quantised-Variational AutoEncoder (VQ-VAE) [61] is a variant of VAE. It introduces the codebook mechanism to quantize the continuous latent representations into discrete components, achieving promising performance on many generative tasks, e.g., text-to-image generation [51], music generation [19, 18], and speech gesture generation [2]. VQ-GAN [21] proposes to promote VQ-VAE by incorporating adversarial loss during the training process. In this paper, we build a 3D-aware VQ-VAE to quantize the triplane features of 3D shapes and acquire discrete geometric parts for autoregressive 3D generation.
@@ 2.3 GPT
Generative Pre-trained Transformer (GPT) [49] originates in the field of natural language processing (NLP). Based on the decoder-only transformer architecture, it autoregressively generates text sequences according to the next-token-prediction paradigm. This series of works [1, 79, 3, 12, 43] with groundbreaking reasoning ability and incredible scalability continue to emerge, revolutionizing language generation. Inspired by these achievements, many researchers have attempted to transfer this scheme to image generation. For example, Parti [72] proposes a pathways autoregressive text-to-image model, which regards image generation as sequence-to-sequence modeling for high-fidelity images. LlamaGen [57] verifies that vanilla autoregressive models, e.g., Llama [60], can achieve state-of-the-art image generation performance without inductive biases on visual signals if scaling properly. Emu3 [65] achieves excellent performance in both generation and perception tasks by tokenizing images, text, and videos into a discrete space. Besides, AutoSDF [46] learns a ‘non-sequential’ autoregressive shape prior for 3D completion, reconstruction, and generation. More recently, several works like MeshGPT [55], MeshAnything [11], and MeshXL [8] also explore to autoregressively generate the faces of 3D meshes with GPT.
@@ 3 Methodology
3D shape encoder is designed to acquire compact and robust latent representations of 3D objects for fine-grained geometry information. Given a 3D object, we first uniformly sample high-resolution point clouds from its surface. To enhance the expressive power of the point clouds, the corresponding normals are also included in the point cloud representation that is denoted as $\mathrm{P}\in\mathbb{R}^{B\times N_{p}\times(3+3)}$, where $B$ is the batch size, and $N_{p}$ is the number of points. Then, we apply the Fourier positional  [58] encoding to the point cloud representation to capture high-frequency details.
Inspired by previous works for point cloud understanding [29, 74], we employ a transformer-based architecture consisting of a cross-attention layer and $N_{1}$ self-attention layers to extract the latent features of the 3D point cloud. More precisely, the point cloud information is injected into a series of learnable query tokens that is denoted as $\mathrm{e}\in\mathbb{R}^{B\times(3\times h\times w)\times d_{e}}$, where $h$ and $w$ are the height and width of the triplane feature maps respectively, and $d$ represents the channel number of these learnable tokens, via the cross-attention layers. After that, the representational ability of these tokens is enhanced by the following self-attention layers, resulting in triplane latent representations, i.e., $\hat{\mathrm{z}}\in\mathbb{R}^{B\times(3\times h\times w)\times d_{z}}$.
@@ 4.1 Experiment Setup
Datasets. To examine the effectiveness of our TAR3D, we conduct experiments on two benchmark 3D datasets, i.e., ShapeNet [6], Objaverse [16], and an out-of-domain dataset, i.e., Google Scanned Object(GSO) [20]. Specifically, the ShapeNet dataset provides 52,472 manufactured meshes covering 55 categories. We follow the splits from 3DILG [73], where 48,597 samples are used for training, 1,283 for validation, and 2,592 for testing. Inspired by previous works on data filtering [33, 76], we score the rendered normal maps of 800,000 meshes in the Objaverse dataset and obtain about 100,000 geometry objects. Moreover, 1,000 samples are randomly selected for performance evaluation, and the remaining ones are employed for model training. In addition, the GSO dataset contains about 1000 real-world 3D scans, which are utilized to further validate the generalization of our method. For each 3D asset in these two datasets, we adopt the rendered images and textual descriptions from ULIP [71] to build the prompt system. We uniformly select 4 images from all the rendered images, whose top 1 captions are utilized as the text prompts.
Metrics. We evaluate the performance of the methods from two aspects: 2D visual quality and 3D geometric quality. For the 2D visual evaluation, we compare the novel views rendered from the synthesized 3D mesh with the ground truth views based on a set of common metrics, including Peak Signal-to-Noise Ratio (PSNR), Perceptual Loss (LPIPS) [77], Structural Similarity (SSIM) [66], and CLIP [50] score. For the 3D geometric evaluation, we compare the point clusters that are randomly sampled from the generated meshes and the ground-truth meshes. Following the protocol in previous works [70, 37], we employ the chamfer distance value and the F-Score with a threshold of 0.02 as the primary evaluation metrics.
Figure 4: Visual comparisons of the 3D meshes generated by our TAR3D and recent multiview-based models, i.e., Syncdreamer, LGM and InstantMesh, and 3D native approaches,i.e., Michelangelo and LN3DIff, for image-to-3D object generation. Given the same input images from the GSO dataset, our TAR3D can produce 3D assets with superior geometric details against other baselines.
@@
In our 3D GPT, we adopt the pre-trained DINO [75](ViT-B16) and FLAN-T5 XL [14] to encode the conditional images and text prompts respectively. As far as the decoder-only transformer, we follow the GPT-L setting of LLamaGen [57], which consists of 24 transformer layers with a head number of 16 and a dimension of 1024. We employ the AdamW optimizer with a learning rate of $41e\mathrm{-}$ 4 and a total batch size of 80 to train our 3D GPT for 100K steps. In addition, the classifier-free guidance (CFG) [25] scale of 7.5 is also introduced in the autoregressive inference to improve geometry quality and image/text-3D alignment.
@@ 4.2 Qualitative Evaluation
Image-to-3D. We first provide visualization comparisons between our TAR3D with recent state-of-the-art models for image-to-3D generation, including three multiview-based methods like SyncDreamer [39], InstantMesh [70], OpenLRM [24], and LGM [59] and three 3D native approaches like Shap-E [30], Michelangelo [78] and LN3Diff [32]. As shown in the garden-swing sample of Fig. 4, the methods based on mulit-view synthesis may produce 3D objects with discontinuous geometry parts or even incorrect objects, which are attributed to inconsistencies in their views used for reconstruction. Meanwhile, the LEGO sample shows that 3D diffusion methods are prone to generating noisy 3D objects, possibly due to anomalies in the conditional denoising process. In contrast to these approaches, our TAR3D model, equipped with the powerful learning capabilities of GPT scheme, can synthesize high-quality 3D objects with superior geometric details.
@@
Methods Shap-E [30] SyncDreamer [39] Michelangelo [78] InstantMesh [70] LGM [59] TAR3D (Ours)
PSNR $\uparrow$ 10.991 11.269 11.928 11.560 11.363 13.626
SSIM $\uparrow$ 0.702 0.706 0.734 0.721 0.714 0.763
Clip-Score $\uparrow$ 0.834 0.837 0.864 0.847 0.841 0.868
LPIPS $\downarrow$ 0.325 0.320 0.278 0.303 0.317 0.216
Chamfer Distance $\downarrow$ 0.156 0.158 0.117 0.137 0.149 0.066
F-Score $\uparrow$ 0.163 0.178 0.226 0.179 0.172 0.303